OpenStudy (anonymous):

24. Solve the equation. 5/13 t = –9 (2 points)

OpenStudy (gw2011):

t=(-9)(13/5) t=-117/5=-23.4

OpenStudy (anonymous):

Multiply both sides by the reciprocal of (5/13) t=(-9)(13/5) t=-117/5

OpenStudy (anonymous):

hey you posted this one already ;)

OpenStudy (anonymous):

i didnt mean to im sry:)

OpenStudy (anonymous):

I don't mind